OpenStudy (anonymous):

Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase has a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius

OpenStudy (anonymous):

SA = 2r(pi) + (pi)r^2 3000 = 2r(pi) + (pi)r^2 (pi)r^2 + 2(pi)r - 3000 = 0 using quadratic formula: r = [-2(pi) +/- SQR((2(pi))^2 - 4(pi)(-3000))]/2(pi) r = [-2(pi) +/- SQR((2(pi))^2 + 12,000(pi))/2(pi) using pi = 3.14: r = [-6.28 +/- SQR(39.4 + 37,680)/6.28 r = [-6.28 +/- 194]/6.28 r = 29.9 or -31.9 using the positive root: r = 29.9cm check: 3000 = 2(29.9)(3.14) + (3.14)(29.9^2) 3000 = 187.8 + 2807.2 3000 = 2995

OpenStudy (anonymous):

I got this far but i didn't put the height in can someone help please

OpenStudy (unklerhaukus):

SA = 2r(pi)h + (pi)r^2

OpenStudy (anonymous):

and where would that go in the equation?

OpenStudy (unklerhaukus):

(pi)r^2 + 2(pi)hr - 3000 = 0

OpenStudy (anonymous):

so at the very end?

OpenStudy (unklerhaukus):

r = [-2(pi)h +/- SQR((2(pi)b)^2 - 4(pi)(-3000))]/2(pi) r = [-2(pi)*50 +/- SQR((2(pi)*50)^2 - 4(pi)(-3000))]/2(pi)

OpenStudy (anonymous):

Can you write the whole equation that i wrote and write in the parts that i forgot

OpenStudy (unklerhaukus):

\[r = \frac{-100\pi \pm \sqrt{(100\pi^2) - 4\pi(-3000)}}{2\pi}\] \[r = \frac{-100\pi \pm \sqrt{10000\pi^2 + 12000\pi}}{2\pi}\]

OpenStudy (anonymous):

do you understand what I am asking?

OpenStudy (unklerhaukus):

not really

OpenStudy (unklerhaukus):

\[r = -50\pm\sqrt{\frac{10000\pi^2}{4\pi^2} + \frac{12000\pi}{4\pi^2}}\]

OpenStudy (unklerhaukus):

\[r = -50\pm\sqrt{2500 + \frac{3000}{\pi}}\]

OpenStudy (anonymous):

I dont have the height yet, i missed it in my problem that I figured out and I just need the height and you to show me where it goes thats it

OpenStudy (unklerhaukus):

well i have put h i n the right places and you see h =50 so the quadratic formula will come out differently

OpenStudy (anonymous):

oh ok well where would it go in my problem

OpenStudy (unklerhaukus):

the lateral surface area of the cylinder is a rectangle one side is the circumference of the circle, the other side in the height of the cylinder, A = 2πr • h

OpenStudy (unklerhaukus):

so the sum of the lateral area and area of one base SA = 2πr•h + πr^2

OpenStudy (anonymous):

ok thank u